solve for x and y - x/a^2 + y/b^2 = 2/ab and x/2b - y/2a = A^3 - B^3/2A^2B^2

Grade:10

1 Answers

Arun
5 years ago
Dear student...this is a very simple question....if you are unable to find answwr then perhaps your calculations has mistake. So please check.However x = a/b while y = b/a